Pan American Airlines’ shares are currently trading at $61.24 each. The market yield on Pan Am's debt is 9% and the firm’s beta is 1.1 The​T-Bill rate is 2% and the expected return on the market is 7.75%. The company’s target capital structure is 70% debt and 30% equity. Pan American Airlines pays a combined federal and state tax rate of 20%. What is the estimated cost of common equity, employing the constant growth dividend discount model? Assume that Pan Am pays annual dividends and that the last dividend of $2.11 per share was paid yesterday. Pan Am started paying dividends 6 years ago. The first dividend was $1.24 per share. 
Answer #1

Dividend Paid 6 years ago = D-6 = $ 1.24 and Current Dividend = D0 = $ 2.11

Average Growth Rate of Dividends = g = [(2.11/1.24)^(1/6)]-1 = 0.09264 or 9.264 %

Current Price = P0 = $ 61.24 and Expected Dividend Next Year = D0 x (1+ Average Growth Rate) = 2.11 x (1.09264) = $ 2.3055 = D1

Therefore, Cost of Common Equity = r = (D1/P0) + g = (2.3055 / 61.24) + 0.09264 = 0.130287 or 13.0287 % ~ 13.03 %